# Online Vectorized HD Map Construction using Geometry

### v1

@@ A.1 Objective Functions
and the simpler one is
$\mathcal{L}^{\prime}=\lambda\cdot\mathcal{L}_{\rm geo}+\beta_{1}\cdot\mathcal{L}_{\rm cls}+\beta_{2}\cdot\mathcal{L}_{\rm pts}+\beta_{3}\cdot\mathcal{L}_{\rm dir}.$ (13)
Point Order Agnostic Matching. In accordance with the methodology proposed by MapTR [18], we employ point order-agnostic matching between predictions and ground truth. In the subsequent formulations, we assume that predictions and ground truth have already been paired.
Classification Loss. To enhance the model’s comprehension of the semantics associated with various map instance types, we incorporate the classification task. Let $\hat{{\bm{p}}}\in\mathbb{R}^{N\times C}$ denote the predicted probabilities, where $C$ is the number of instance categories. Here, $\hat{{\bm{p}}}_{ic}$ represents the predicted probability of instance $i$ belonging to category $c$. With ground truth labels ${\bm{y}}\in\{1,...,C\}^{N}$, the objective function based on focal loss is defined as follows:
$\mathcal{L}_{\rm cls}=-\sum_{i=1}^{N}\sum_{c=1}^{C}\delta[{\bm{y}}_{i}=c]\cdot\alpha_{c}(1-\hat{{\bm{p}}}_{ic})^{\gamma}\log\hat{{\bm{p}}}_{ic}$ (14)
@@
Point Loss. For the perception of instance positions, we employ a point loss that evaluates $L^{2}$ distances between predicted points and ground truth points, which is specified as
$\mathcal{L}_{\rm pts}=\sum_{i=1}^{N}\sum_{j=1}^{N_{v}}\|\hat{{\bm{L}}}_{j}^{i}-{\bm{L}}_{j}^{i}\|_{2}.$ (15)
Edge Direction Loss. To obtain more precise displacement vectors, which are crucial in our G-Representation, we incorporate an edge direction loss. This loss quantifies the cosine similarity between predicted displacement vectors and their corresponding ground truth vectors. Specifically, the loss is defined as
@@
Segmentation Loss. The auxiliary binary segmentation task is valuable for assisting the model in the coarse perception of shape geometry. We integrate a convolutional neural network-based BEV segmentation head with BEV features. Let $\hat{{\bm{P}}}_{\rm bev}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ represent the probability of each grid belonging to the instance area, and ${\bm{Y}}_{\rm bev}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$ denote the ground truth. The corresponding objective function is defined as
$\mathcal{L}_{\rm bev}=\mathcal{L}_{\rm bce}(\hat{{\bm{P}}}_{\rm bev},{\bm{Y}}_{\rm bev})$ (17)
where the binary cross entropy loss $\mathcal{L}_{\rm bce}$ is
@@
We also introduce the auxiliary PV segmentation task, incorporating a shared convolutional neural network head for all views. The ground truth is projected back to the PV space to form the binary mask. Let $\hat{{\bm{P}}}_{\rm pv}^{k}\in\mathbb{R}^{H\times W}$ denote the segmentation results for view $k$ with corresponding ground truth ${\bm{Y}}_{\rm pv}^{k}\in\{0,1\}^{H\times W}$, then the objective function can be expressed as
@@
Finally, we obtain the segmentation loss as
@@
Depth Estimation Loss. To enhance depth perception, we replace GKT [2] with LSS [30] when applying the full objective function. Let $\hat{{\bm{P}}}^{k}_{\rm dep}\in\mathbb{R}^{H\times W\times D}$ represent the depth distribution of each grid estimated by LSS in the PV space of view $k$, where $D$ represents the number of quantified depth buckets. Given the ground truth ${\bm{Y}}_{\rm dep}^{k}\in\{1,...,D\}^{H\times W\times D}$, the depth estimation loss is defined as
@@ A.2 Hyperparameter Settings
Appendix B More Experimental Results
In this section, we present additional ablation studies and hyperparameter experiment results. In all of these experiments, the model is trained for 24 epochs on NuScenes using the simplified objective function. Unless otherwise specified, we employ the default settings outlined in § A.2.
B.1 Decoupled Attention
In the proposed GDA, we decouple shape and relative perception, sequentially applying Euclidean shape and relation attention. For further ablation, we also experiment with naively doubling self-attention layers in one block, denoted as 2-SA. As shown in Table A3, the results indicate the superiority of GDA when cooperating with Euclidean Loss. This outcome is also suggested by the results in § B.2, as naively increasing the number of attention blocks does not lead to an improvement in the model’s performance. The decoupling of the processing of shape and relation geometry plays a crucial role in our design.
Table A3: GDA collaborates more effectively with Euclidean loss than two layers of vanilla self-attention.
[…]
2-SA 53.5 46.2 54.4 51.4
GDA 53.6 49.6 54.7 52.6
B.2 Impact of the Decoder Block Number
@@
[…]
@@
B.3 Impact of the Query Number
@@
Instance Queries. As depicted in Table A5, augmenting the number of instance queries could be advantageous for the model’s performance. More specifically, the mAP exhibits an increment of $+2.2\%$ when the query number is elevated from 50 to 100. This observation aligns with intuition, as a higher number of instance queries implies a broader pool of diverse candidates. Though a larger number of instance queries might be beneficial, we set $N=50$ in our major experiments to align with the parameters employed in prior studies [18, 31], ensuring a fairer comparison.
@@
[…]
@@
[…]
@@
Appendix C More Visualization Results
We present additional visualization cases under varied weather conditions, as illustrated in Figure A1 to Figure A3. Our method is trained with a ResNet-50 backbone using the simplified objective function.
As illustrated in Figure A1, even in challenging rainy conditions, our method demonstrates more robust results. Particularly in scenario (d) of Figure A1, where the highlighted road boundary and lane divider are heavily occluded by water on the front windshield, our method can still recover the entire instance from observed parts. This showcases the potential of the proposed geometric designs.

### v2

@@ 0.A.1 Objective Functions
and the simpler one which excludes dense prediction losses is:
$\mathcal{L}^{\prime}=\lambda\cdot\mathcal{L}_{\rm Euc}+\beta_{1}\cdot\mathcal{L}_{\rm cls}+\beta_{2}\cdot\mathcal{L}_{\rm pts}+\beta_{3}\cdot\mathcal{L}_{\rm dir}.$ (12)
Point Order Agnostic Matching. In accordance with the methodology proposed by MapTR [20], we employ point order-agnostic matching between the prediction and ground truth. In the subsequent formulations, we assume that the prediction and ground truth have already been paired.
Classification Loss. To enhance the model’s comprehension of semantics associated with various map instance types, we incorporate the classification task. Let $\hat{{\bm{p}}}\in\mathbb{R}^{N\times C}$ denote the predicted probabilities, where $C$ is the number of instance categories. Here, $\hat{{\bm{p}}}_{ic}$ represents the predicted probability of instance $i$ belonging to category $c$. With ground truth labels ${\bm{y}}\in\{1,...,C\}^{N}$, the objective function based on focal loss is defined as follows:
$\mathcal{L}_{\rm cls}=-\sum_{i=1}^{N}\sum_{c=1}^{C}\delta[{\bm{y}}_{i}=c]\cdot\alpha_{c}(1-\hat{{\bm{p}}}_{ic})^{\gamma}\log\hat{{\bm{p}}}_{ic}\,,$ (13)
@@
Point Loss. For the perception of instance positions, we employ a point loss that evaluates $L^{1}$ distances between predicted points and ground truth points, which is specified as:
$\mathcal{L}_{\rm pts}=\sum_{i=1}^{N}\sum_{j=1}^{N_{v}}\|\hat{{\bm{L}}}_{j}^{i}-{\bm{L}}_{j}^{i}\|_{1}.$ (14)
Edge Direction Loss. To obtain more precise displacement vectors, which are crucial in our G-Representation, we incorporate an edge direction loss. This loss quantifies the cosine similarity between predicted displacement vectors and their corresponding ground truth vectors. Specifically, the loss is defined as:
@@
Segmentation Loss. The auxiliary binary segmentation task is valuable for assisting the model in the coarse perception of shape geometry. We integrate a convolutional neural network-based BEV segmentation head with BEV features. Let $\hat{{\bm{P}}}_{\rm bev}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ represent the probability of each grid belonging to the instance area, and ${\bm{Y}}_{\rm bev}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$ denote the ground truth. The corresponding objective function is defined as:
$\mathcal{L}_{\rm bev}=\mathcal{L}_{\rm bce}(\hat{{\bm{P}}}_{\rm bev},{\bm{Y}}_{\rm bev})\,,$ (16)
where the binary cross entropy loss $\mathcal{L}_{\rm bce}$ is:
@@
We also introduce the auxiliary PV segmentation task, incorporating a shared convolutional neural network head for all views. The ground truth is projected back to the PV space to form the binary mask. Let $\hat{{\bm{P}}}_{\rm pv}^{k}\in\mathbb{R}^{H\times W}$ denote the segmentation results for view $k$ with corresponding ground truth ${\bm{Y}}_{\rm pv}^{k}\in\{0,1\}^{H\times W}$, then the objective function can be expressed as:
@@
Finally, we obtain the segmentation loss as follows:
@@
Depth Estimation Loss. To enhance depth perception, we adopt an auxiliary depth estimation task. Let $\hat{{\bm{P}}}^{k}_{\rm dep}\in\mathbb{R}^{H\times W\times D}$ represent the depth distribution of each grid estimated by LSS [33] in the PV space of view $k$, where $D$ represents the number of quantified depth buckets. Given the ground truth ${\bm{Y}}_{\rm dep}^{k}\in\{1,...,D\}^{H\times W\times D}$, the depth estimation loss is defined as:
@@ 0.A.2 Hyperparameter Settings
Appendix 0.B More Experimental Results
In this section, we present additional ablation studies and hyperparameter experiment results. In all of these experiments, the model is trained for 24 epochs on nuScenes using the simplified objective function. Unless otherwise specified, we employ the default settings outlined in § 0.A.2.
0.B.1 Impact of the Decoder Block Number
@@
# Block APdiv()($\uparrow$)( ↑ ) APped()($\uparrow$)( ↑ ) APbnd()($\uparrow$)( ↑ ) mAP()($\uparrow$)( ↑ )
@@
0.B.2 Impact of the Query Number
@@
Instance Queries. As depicted in Table A4, augmenting the number of instance queries could be advantageous for the model’s performance. More specifically, the mAP exhibits an increment of $+27.8\%$ when the query number is elevated from 10 to 50. This observation aligns with intuition, as a higher number of instance queries implies a broader pool of diverse candidates.
@@
$N$ APdiv()($\uparrow$)( ↑ ) APped()($\uparrow$)( ↑ ) APbnd()($\uparrow$)( ↑ ) mAP()($\uparrow$)( ↑ )
@@
$N_{v}$ APdiv()($\uparrow$)( ↑ ) APped()($\uparrow$)( ↑ ) APbnd()($\uparrow$)( ↑ ) mAP()($\uparrow$)( ↑ )
@@
Appendix 0.C More Visualization Results
We present additional visualization cases under varied weather conditions, as illustrated in Figure A1 to Figure A3. Our method is trained with a ResNet50 backbone using the simplified objective function.
As illustrated in Figure A1, in challenging rainy conditions, our method demonstrates more robust results. Particularly in scenario (d) of Figure A1, where the front road boundary and lane divider are heavily occluded by water on the front windshield, our method can still recover the entire instance accurately from observed parts. This showcases the potential of proposed geometric designs.